### v1

# Controllable Prompt Tuning For

Abstract
Models trained on data composed of different groups or domains can suffer from severe performance degradation under distribution shifts. While recent methods have largely focused on optimizing the worst-group objective, this often comes at the expense of good performance on other groups. To address this problem, we introduce an optimization scheme to achieve good performance across groups and find a good solution for all without severely sacrificing performance on any of them. However, directly applying such optimization involves updating the parameters of the entire network, making it both computationally expensive and challenging. Thus, we introduce Controllable Prompt Tuning (CPT), which couples our approach with prompt-tuning techniques. On spurious correlation benchmarks, our procedures achieve state-of-the-art results across both transformer and non-transformer architectures, as well as unimodal and multimodal data, while requiring only $0.4\%$ tunable parameters.
1 Introduction
Spurious correlation or shortcut learning arises when a classifier relies on non-predictive features that coincidentally correlate with class labels among training samples. For example, a majority of waterbirds are often found near water, while landbirds mostly appear near land. Even when being provided with a small number of minority group examples, such as waterbirds on land, or landbirds on water, model predictions are often still dominated by spurious features, in this case, the background, ignoring the salient feature, the foreground [68].
A variety of approaches aim to improve worst group performance, by oversampling high-loss samples [53, 90], undersampling the majority group [69] or re-training a last layer on group-balanced validation data [42]. A particularly popular and effective baseline, Group Distributionally Robust Optimization (GroupDRO) [68, 69], directly minimizes an estimated upper bound on worst group loss. However, all these procedures generally neglect knowledge transfer between groups. Moreover, high-error samples can potentially include noisy-label data that could hurt the model’s predictive ability if we try to learn them exhaustively [61]. Besides, DRO-based methods, in particular, are susceptible to overfitting, wherein their test performances can decline to the same as ERM with sufficiently long training [31, 64, 88].
Figure 1: Accuracy curves on Waterbirds for four groups during training. Vertical lines indicate early stopping epochs when models perform best on the validation set.
To visually illustrate these limitations, we plot in Figure 1 the performance of a ResNet50 on the Waterbirds benchmark, trained with ERM, GroupDRO, and our approach (to be explained later). The plots demonstrate that while GroupDRO can improve the performance of the minority group (green line) early in training compared to ERM, it rapidly overfits and fails to maintain this performance over time. After training for ten epochs, the test performance gap between the minority and majority (blue line) groups grows sharply. To address these critical limitations, we introduce a systematic approach to train a model that exhibits consistently high performance across all groups. Specifically, at each iteration, we identify a descending direction that benefits the groups simultaneously rather than focusing on an individual group. Figure 1 shows that our proposed method performs almost equally well in every group, thanks to our balancing mechanism.
Moreover, we aim to control the group-wise learning process by ensuring that the magnitude of loss for different groups is inversely proportional to a predefined vector $\mathbf{c}$. Varying this hyperparameter $\mathbf{c}$ adjusts the trade-off between the worst group and the average performance, yielding a more flexible method than prior work that only focuses on maximizing the worst group accuracy. However, computation can grow linearly with the number of controlling vectors $\mathbf{c}$. To overcome this challenge, we propose decoupling our optimization framework with parameter-efficient fine-tuning (PEFT) techniques that considerably reduce the number of trainable parameters. With the model compression ability powered by prompt-tuning, our proposed model can scale well for a large number of controlling vectors with a small parameter increase, e.g., $0.01\%$ and $0.003\%$ per value of $\mathbf{c}$ for Vision Transformer [23] and CLIP [66], respectively.
Contribution. In this work, we introduce CPT, a Controllable Prompt Tuning method to prevent models from learning spurious features. In summary, our key contributions are depicted as follows:
  • Drawing inspiration from the principles of multi-objective optimization theory, we introduce a novel balancing mechanism to address the challenges of group distributional robustness. Our method not only considers the worst group but also leverages the gradient information from all groups to determine an ultimate updating direction that benefits all of them.
  • While the above rigid balancing procedure works well in different scenarios, we extend our method by introducing a controlling vector that allows us to dynamically adjust the priority across groups. Additionally, we integrate prompt tuning techniques to enhance the scalability of our method as the complexity and number of controlling vectors increase and to make it applicable for de-biasing large transformers-based models.
  • We conduct intensive experiments on many different benchmarks, where CPT consistently exhibits superior performance. Our method surpasses the current state-of-the-art baselines on Waterbirds and CelebA datasets while updating $0.4\%$ parameters. Moreover, it also outperforms recently proposed methods that aim to de-bias Vision Transformer and CLIP models with minimal training cost.
2 Related work
We first review relevant methods to enhance distributional robustness and then discuss prior work related to the two essential parts of our proposed method: transformer and parameter-efficient fine-tuning.
Distributional robustness
can be compromised for reasons like selection bias or spurious correlation across tasks. For selection bias on individual samples, prior works seek to reduce bias via practices like importance reweighting [36, 72, 16, 46], hard sample reweighting [53, 60] or distribution matching and discrepancy minimization [17, 7, 8, 63], and domain-adversarial algorithms [27, 58] across tasks in their feature representation space. For selection bias on groups or subpopulation, namely subpopulation shift or dataset imbalance, label propagation  [11, 8] or other consistency regularization [59, 84] are used to generalize the prediction to broader domains. In particular, GroupDRO [68, 69] or other relevant methods [44, 39] select descending direction more in favor of the worse performing/minority group(s) during training.
Improving distributional robustness also requires mitigating the effect of environmental features. Prior work uses model ensemble [45] or model soups [81] to learn rich features and reduce the effect of spurious features. More practices include invariant feature representation learning [6, 14], domain generalization [44, 73, 71] or through data augmentation [83, 86].
Transformer In the past few years, transformer [77] has emerged as a dominant architecture for a wide range of applications, exhibiting their profound capacity in modeling different modalities: languages [21, 55, 76], vision [23, 12], speech [67, 79], and time series [49, 30]. Due to its ability to process multiple modalities, transformer-based architectures were widely adopted for image understanding using different types of supervision. For example, Vision Transformer (ViT) [23] utilizes masked patches for pre-training and image-label supervision for fine-tuning. Notably, Contrastive Language-Image Pre-training (CLIP) [66], which was pre-trained on billion-scale image-text pairs, has demonstrated excellent zero-shot performance in image classification tasks. Despite obtaining impressive results on various downstream tasks, over-parameterization is still a crucial problem for transformer [26, 62] which potentially causes overfitting [48] and unintended biases [1, 80, 24, 89] more severely than small-scale architectures [69].
Parameter-Efficient Fine-Tuning While conventional training updates the entire network, prior work has shown that it is possible to achieve performance comparable with full fine-tuning by instead updating a small set of parameters [51, 33, 34]. Early work in this line of research includes adapter [37, 13, 74], which adds a lightweight fully connected network between the layers of a frozen pre-trained model. Since then, there have been many other methods that update subsets of the network [32, 28, 87] or reparametrize the weights of the network using low-rank decompositions [2, 38, 25]. Another appealing PEFT strategy is prompt tuning, which prepends a few trainable tokens before the input sequence or hidden state to encode task-specific knowledge to the pre-trained model [47, 50, 94, 40].
3 Background
In this section, we first formulate the problem of group robustness and recap the technical details of GroupDRO, then revisit some concepts of multi-objective optimization.
3.1 Problem formulation
Formally, we consider the conventional setting of classifying an input $x\in\mathcal{X}$ as a target $y\in\mathcal{Y}$, where $\mathcal{X},\mathcal{Y}$ are input and label spaces, respectively. We are given a training dataset composed of $K$ groups from the set $\mathcal{G}$, where each group $g\in\mathcal{G}$ consists of $n_{g}$ instances sampled from the probability distribution $P_{g}(\mathcal{X},\mathcal{Y})$. Since the numbers of examples are different among groups in the training data, we consider groups with relatively large $n_{g}$ as majority groups and those with small $n_{g}$ as minority groups. Our main goal is to develop a model that performs effectively across all groups within $\mathcal{G}$ to avoid learning spurious correlations.
Following previous methods [69], we adopt two metrics: worst group accuracy, indicating the minimum test accuracy across all groups $g\in\mathcal{G}$, and average accuracy, which represents the weighted average test accuracy with the weights corresponding to the relative proportions of each group in the training data. It is worth noting that commonly used datasets for group robustness often exhibit skewed training set distributions, the weighted average score is thus dominated by the performance of the model on groups that include large numbers of training data. Therefore, we propose to evaluate the model performance by additionally reporting the mean (unweighted average) accuracy score across different groups.
3.2 Group distributionally robust optimization
GroupDRO learns the model parameters $\theta$ by directly optimizing for the worst-group training loss as follows:
$\max_{g\in\mathcal{G}}\mathbb{E}_{(x,y)\sim P_{g}(\mathcal{X},\mathcal{Y})}\left[\ell\left(f_{\theta}(x),y\right)\right]$
for some loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$. Then, their updating formula at each step is given by:
$\quad g^{*}=\arg\max\ell_{g}(\theta)\quad$ and $\quad\theta^{t+1}=\theta^{t}-\eta\nabla_{\theta}\ell_{g^{*}}$,      where $\eta$ is the learning rate.
3.3 Multi-objective Optimization
Assume that we are given $m$ objectives functions $f_{i}(x)$, $i\in[m]$ where $x\in\mathbb{R}^{d}$ and each $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a scalar-valued function (we use the $[m]$ notation to denote the set $\{1,2,\dots,m\}$) . In the multi-objective optimization (MOO) problem, we are interested in minimizing a vector-valued objective function whose $i$-th component is $f_{i}(x)$:
$\vec{F}(x)=[f_{1}(x),f_{2}(x),\dots,f_{m}(x)]$ (1)
In general, an optimal solution for this objective vector function does not exist since each of the individual function $f_{i}(x)$ does not necessarily guarantee to have the same minimum solution. Alternatively, we expect to obtain a solution, from which we cannot improve any specific objective without hurting another. According to the above formulation, the solution [95] of the multi-objective minimization problem is formally defined as follows:
Definition 1.
(Pareto dominance) Let $x_{1},x_{2}$ be two solutions for the multi-objective optimization problem in Equation 1, $x_{1}$ dominates $x_{2}$ $(x_{1}\prec x_{2})$ if and only if $f_{i}(x_{1})\leq f_{i}(x_{2})\forall i\in[m]$ and $\exists j\in[m]$ s.t. $f_{j}(x_{1})<f_{j}(x_{2})$.
Definition 2.
(Pareto optimality) A solution $x^{*}$ in problem 1 is said to be Pareto optimal if it is not dominated by any other solution. Therefore, the solutions set of the multi-objective minimization problem is given by $\mathcal{P}:=\left\{x\mid\nexists x^{\prime}:x^{\prime}\prec x\right\}$. This set of all Pareto optimal solutions is called the Pareto set while the collection of their images in the objective space is called the Pareto front.
Gradient-based MOO Désidéri [20] shows that the descent direction can be found in the convex hull composed of gradients corresponding to different objectives $\{\boldsymbol{h}^{i}:=\nabla f_{i}\}_{i=1}^{m}$:
$\mathcal{CH}=\left\{\boldsymbol{H}^{\top}\boldsymbol{\alpha}=\sum_{i=1}^{m}\alpha_{i}\boldsymbol{h}^{i}\mid\alpha\in\Delta_{m}\right\}$
where $\Delta_{m}$ denotes the $m-1$ dimension simplex.
Moreover, they introduce the Multiple Gradient Descent Algorithm (MGDA), which calculates the minimum-norm gradient vector $\boldsymbol{h}$ that lies in the convex hull: $\boldsymbol{h}=\text{argmin}_{h\in\mathcal{CH}}||h||_{2}$. This approach can guarantee that the obtained solutions lie on the Pareto front, from which we cannot find an updating direction that decreases all objectives simultaneously.
4 Proposed method
In GroupDRO, the updating formula solely focuses on the worst-performing group, which is susceptible to challenges when groups exhibit varying loss magnitude, noise, and transfer characteristics. For example, among the high-loss group, there might exist a few outliers that have large input shifts or even wrong labels, which can raise the training loss of the whole group and prevent the model from fitting other groups.
4.1 Balancing group distributional robustness
Figure 2: Overview of our method on the Waterbirds dataset. Our main objective is not only to improve model performance across groups by optimizing their loss functions $\ell_{1},\ell_{2},\ell_{3},\ell_{4}$, but also to maximize the entropy over this loss distribution.
In this paper, we go beyond the worst group approach of GroupDRO to examine all groups at a time. To encourage the model to learn from different groups at each step to avoid struggling on just the hardest one only, we propose to minimize the following $K$-dimension loss vector:
$\vec{\mathcal{L}}(\theta)=\big{\{}\mathbb{E}_{(x,y)\sim P_{g^{k}}(\mathcal{X},\mathcal{Y})}\left[\ell\left(f_{\theta}(x),y\right)\right]\big{\}}_{k=1}^{K}$ (2)
where $K$ is the number of groups and $g^{k}$ is the $k$-th group.
At $t$-th iteration, we search for an update vector that minimizes the weighted sum of group losses:
$d_{t}(w)=w^{\top}\nabla\vec{\mathcal{L}}(\theta_{t})=\sum_{k=1}^{K}w_{k}\nabla\ell_{k}(\theta_{t})$ (3)
Then, the updating formula of our algorithm is then given by $\theta_{t+1}=\theta_{t}-\eta d_{t}(w)$ for some step size $\eta\in\mathbb{R}^{+}$. The coefficient $w$ is chosen such that our update rule can benefit all groups $d_{t}(w)^{\top}\nabla\ell_{k}(\theta_{t})\geq 0$. To this end, a straightforward approach that is readily applicable to optimize the objective (2) is applying MGDA to find the solution of (3) that yields the minimum Euclidean norm of the composite gradient:
$d_{t}^{*}=\operatorname*{arg\,min}_{w\in\Delta_{{}_{K}}}\left\|\sum_{k=1}^{K}w_{k}\nabla\ell_{k}(\theta_{t})\right\|_{2}$ (4)
However, MGDA is often biased toward objectives with smaller gradient magnitudes [54], and lacks controllable property (i.e. could not adjust the group that we want to focus more or less at each update), which can be utilized to balance the loss between groups. Hence, we introduce a new objective that steers the updating direction to obtain our desired balanced magnitude among groups. Assuming that the entropy $H(p)$ measures the diversity of a continuous distribution $p$, this behavior could be achieved via maximization of the entropy with respect to the distribution of loss functions:
$\mathcal{L}_{ent}(\theta)$ $=H\big{(}\text{softmax}\big{(}\vec{\mathcal{L}}(\theta))\big{)}=H\Big{(}\text{softmax}\big{(}\big{\{}\mathbb{E}_{P_{g^{k}}}\left[\ell\left(f_{\theta}(x),y\right)\right]\big{\}}_{k=1}^{K}\big{)}\Big{)}$ (5)
Theorem 3.
Assume that the loss function $\ell$ is differentiable up to the first order with respect to $\theta$, then following
$d_{ent}:=\sum_{i=1}^{K}\nabla\ell_{i}(\theta)\Big{[}p_{i}\log(p_{i})-p_{i}\sum_{j=1}^{K}\log(p_{j})p_{j}]\Big{]}$
where $p_{i}=\frac{e^{\ell_{i}(\theta)}}{\sum_{j=1}^{K}e^{\ell_{j}(\theta)}}$, maximizes the objective $\mathcal{L}_{ent}(\theta)$.
The proof of Theorem 3 is deterred to Appendix A. Theorem 3 provides us with the updating rule that increases the entropy term in Equation (5) the most. To balance the learning across groups, we seek the coefficient $w$ that maximizes the cosine similarity with $d_{ent}$ and does not conflict against the gradient vector of each group. Denotes $w^{\prime}_{i}=p_{i}\log(p_{i})-p_{i}\sum_{j=1}^{K}\log(p_{j})p_{j}$, this is equivalent to optimizing:
$w^{*}$ $=\operatorname*{arg\,max}_{w\in\Delta_{K}}d_{t}(w)^{\top}d_{ent}$ (6)
[…]
While the balance between loss magnitude is not achieved, we strictly follow the updating direction that decreases training errors of high-loss groups but slightly relaxes constraints in the optimization problem (6) to allow bounded increases among groups that already have small loss magnitudes.
$w^{*}=\operatorname*{arg\,max}_{w\in\Delta_{K}}d_{t}(w)^{\top}d_{ent}$
$\;\;\;\;=\operatorname*{arg\,max}_{w\in\Delta_{K}}w^{\top}\nabla\vec{\mathcal{L}}(\theta)^{\top}\nabla\vec{\mathcal{L}}(\theta)w^{\prime}$ (7)
[…]
$\;\;d_{t}(w)^{\top}\nabla\ell_{k}(\theta_{t})\geq 0\;\text{if}\;\ell_{k}(\theta_{t})=\max_{k\in[K]}\ell_{k}(\theta_{t}).$
$\;\;d_{t}(w)^{\top}\nabla\ell_{k}(\theta_{t})\geq d_{ent}^{\top}\nabla\ell_{k}(\theta_{t})\;\text{if}\;\ell_{k}(\theta_{t})<\max_{k\in[K]}\ell_{k}(\theta_{t}).$
The second constraint guarantees that the obtained solution is not worse than $d_{ent}$ on those groups in which the model is performing well or when $d_{ent}$ itself intrinsically supports learning them. Once the loss magnitude is almost balanced ($||d_{ent}||\leq\epsilon$), we simply update the model parameter by the average group gradients, i.e. $w=(\frac{1}{K},\frac{1}{K},\dots,\frac{1}{K})$.
In summary, the key difference between our proposed method and GroupDRO is that we solve a small $K$-dimensional optimization problem to find the updating direction that not only focuses on the worst group but also improves the performance in other groups and balances the loss magnitude between them at the same time. In practical settings, the group number $K$ is often a small number (e.g. $K=4$), thus could be efficiently solved using standard linear programming libraries [22, 5] without introducing significant computational overhead (Appendix B).
4.2 Controllable prompt tuning for balancing group distributional robustness
To achieve the controllable property, we apply the optimization procedure in section 4.1 on $\ell_{k}(\theta)={c}_{k}\mathbb{E}_{P_{g^{k}}}\left[\ell\left(f_{\theta}(x),y\right)\right]$ where $\mathbf{c}\in\mathbb{R}^{K}$ is a pre-defined vector that is used to adjust the magnitude of group losses. Intuitively, the principle behind our proposed objective is to enforce the loss vector $\vec{\mathcal{L}}(\theta)$ to be inversely proportional to $\mathbf{c}$, from which we can adjust the group loss magnitude, and thus control the trade-off between them by varying $\mathbf{c}$. For example, we can put more weight on penalizing loss terms of some specific groups to accelerate the learning progress on those groups, or conversely slow down the learning to avoid overfitting on oversampled groups [19].
In practice, tuning the value of $\mathbf{c}$ to obtain our desired model behavior requires learning multiple models individually and independently. Therefore, the computational complexity of our proposed method can grow linearly with the number of vectors $\mathbf{c}$, thus being very expensive in the context of deep learning. To reduce overhead, we adopt the usual practice of parameter-efficient fine-tuning methods and optimize only a small portion of the model. In particular, we freeze the main backbone while introducing lightweight prompt sets, one for each task. For example, we employ the same ViT backbone for the two datasets Waterbirds, CelebA, but using different sets of prompts (Section 5).
Overview of our proposed method is depicted in Figure 2, given an input image belonging to one of a group $\{g^{1},g^{2},\dots,g^{K}\}$ and a set of classes, we solve the optimization problem (4.1) to find the reweighting coefficient $w$ that maximizes the entropy $\mathcal{L}_{ent}(\theta)$ while reasonably decreases group losses. While Figure 2 illustrates our proposed method on the CLIP model on the image classification task, it applies to other transformer-based models and task types. In our experiment, we exploit both image-end prompts and language-end prompts tuning on frozen ViT and CLIP backbones, respectively.
Prompt tuning for Vision Transformer: Regarding the prompt design for a ViT of $N$ layers, we use the set of continuous learnable tokens $\{\mathbf{P}_{n}\}_{n=0}^{N-1}$. Each $\mathbf{P}_{n}\in\mathbb{R}^{L\times D}$, where $D$, $L\in\mathbb{N}$ indicate the prompt length and latent space dimension, respectively. Subsequently, the input at the $(n+1)$-th layer follows the format of $\left[x_{n},\mathbf{P}_{n},\mathbf{E}_{n}\right]$, where $\mathbf{x}_{n},\mathbf{E}_{n}\in\mathbb{R}^{D}$ represent the [CLASS] token and image patch embeddings after $n$-th layer while [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operator. Thus, the number of trainable parameters for this image-end prompt set is $L\times N\times D$.
Prompt tuning for CLIP: Being trained on an enormous amount of image-text pairs, CLIP offers an expressive representation for both language and image inputs by using two separate encoders, one for each modality, as shown in Figure 2. Since the image encoder can be either ResNet [35], ConvNeXt [57] or ViT [23], we conduct prompt tuning on the transformer text encoder for consistency. In short, we employ a learnable context prompt of length $L$: $\mathbf{P}\in\mathbb{R}^{L\times D}$, to replace the hand-crafted prompt used by the original CLIP (i.e., “this is a photo of a "). Based on this design, the input for the text encoder of each class has the format of $\left[x,\mathbf{P}\right]$, where $x$ is the [CLASS] embedding of the corresponding class name and has the same dimension of $D$. These learnable tokens $\mathbf{P}$ are better at capturing domains-specific knowledge than those artificial prompt [92, 93] and can be learned to de-bias the original CLIP model while introducing a few trainable parameters ($N\times D$).
5 Experiments
In this section, we evaluate the effectiveness of the proposed CPT method on benchmark image datasets in the presence of spurious features: Waterbirds [68], CelebA [56], MetaShift [52] and ISIC [15]. Due to the space limit constraint, we briefly provide an overview of our experimental setup below, see Appendix B for more details and additional results.
5.1 Datasets
Waterbirds is created by placing bird photos from the Caltech-UCSD Birds dataset [78] with background images taken from Places [91]. The target attributes are foreground objects (i.e. landbird or waterbird) while spurious correlations are background (i.e. land or water landscape).
[…]
MetaShift is another real-world dataset [52] for cat-dog classification, where the cat and dog images are spuriously correlated with indoor (e.g. bed) and outdoor objects (e.g. bike). For evaluation, models are given dog and cat images associated with shelf background, which does not appear in the training set. Thus, this dataset represents a more coherent distributional shift, compared to other benchmarks.
ISIC is an even more challenging dataset with the appearance of multiple spurious features [15], which includes dermoscopic images of skin lesions with target attributes benign or melanoma. Visual examples of those datasets, along with the size for each group in train, validation, and test sets are given in Appendix B.
5.2 Baselines
Following the previous studies, we compare CPT against other state-of-the-art methods for mitigating spurious correlations, including GroupDRO [68], Deep Feature Reweighting (DFR) [42], Progressive Data Expansion (PDE) [19], Subsample [19], Common Gradient Descent (CGD) [64] with the access to the group information during training. Besides, we consider those methods require group labels for validation: Learning from failure (LfF) [60], Just Train Twice (JTT) [53], Predictive Group Invariance (PGI) [3], Contrastive Input Morphing (CIM) [75], Correct-n-Contrast (CNC) [90], or those do not use group annotations at all like Empirical Risk Minimization (ERM) and Automatic Feature Reweighting (AFR) [65]. Moreover, we also report the results from two recent studies of spurious correlation on ViT [29] and CLIP [85].
Regarding the baselines for those experiments further taking the distribution shift into account, we follow exactly the protocol of [82] and compare our method to Empirical Risk Minimization (ERM), Upweighting (UW), Invariant risk minimization (IRM) [6], Information Bottleneck based Invariant Risk Minimization (IB-IRM) [4], Variance Risk Extrapolation (V-REx) [44], CORrelation ALignment (CORAL) [73], and Fish [71]; instance reweighting methods: GroupDRO [68], JTT [53], Domain mixup (DM-ADA) [83], Selective Augmentation (LISA) [86] and the current state-of-the-art method on those benchmarks Discover and Cure (DISC) [82].
5.3 Model training and evaluation
We examine our proposed method on ResNet50 [35], ViT-B/16 [23] and different variants of OpenAI’s CLIP [66] to align with previous work [42, 29, 85]. For prompt tuning on ViT, we employ prompts of lengths $L=5$ in the MetaShift and ISIC, or $L=10$ in the Waterbirds and CelebA datasets. Similarly, $L$ is set to $16$ for all CLIP backbones throughout our experiment. Unless otherwise explicitly mentioned in the experiment, the value of the controlling vector $c$ is set to the default value of $(1,\dots,1)$. Other detailed configurations are relegated to Appendix B.
Due to the large number of groups in ISIC, we calculate the AUROC score [10], following [82, 9] while computing average, worst-group and mean (if possible) accuracy scores for other datasets. For a fair and precise comparison, results of baselines are taken from recent studies [19, 82, 64, 90, 41] and reported directly from original papers.
5.4 Results
Waterbirds and CelebA We start by validating CPT on two well-established benchmarks for investigating spurious correlation, namely Waterbirds and CelebA in Table 5.4. From the results, we see that CPT consistently achieves the best worst-group accuracy compared with the recent state-of-the-art methods such as DFR and PDE on both datasets. Remarkably, our method only updates a tiny portion ($0.4\%$) of parameters. As such, CPT is much more computationally efficient than methods that need to update the entire network. Furthermore, we see that CPT has strong gains in average accuracy, thus not only outperforming other spurious correlation methods on both metrics but also closing the gap to the average accuracy of ERM. Another interesting observation is that those baselines aim to balance the learning between groups via downsampling the minority groups (Subsample), reweighting group losses inversely proportional to the group populations (UW), or computing scaling weights using group training performance to upweight the worst-group loss (GroupDRO), are effective in alleviating spurious correlation. Even though, CPT still exhibits impressive performance compared to the aforementioned methods by using an adaptive balancing mechanism at each iteration instead of relying on fixed reweighting coefficients.
Table 1: Overall results on Waterbirds and CelebA datasets with best methods are highlighted in bold. Performance is evaluated on the test set with models early stopped at the highest worst-group accuracy on the validation set.
Train group info Validation group info Train once Waterbirds CelebA
Method Worst Average # Params Worst Average # Params
ERM $\times$ $\times$ $\surd$ $70.0_{\pm 2.3}$ $97.1_{\pm 0.1}$ $23$M $45.0_{\pm 1.5}$ $94.8_{\pm 0.2}$ $23$M
AFR $\times$ $\times$ $\times$ $90.4_{\pm 1.1}$ $94.2_{\pm 1.2}$ $23$M $82.0_{\pm 0.5}$ $91.3_{\pm 0.3}$ $23$M
LfF $\times$ $\surd$ $\times$ $78.0_{N/A}$ $91.2_{N/A}$ $23$M $77.2_{N/A}$ $85.1_{N/A}$ $23$M
SSA $\times$ $\surd$ $\times$ $89.0_{\pm 0.6}$ $92.2_{\pm 0.9}$ $23$M $89.8_{\pm 1.3}$ $92.8_{\pm 1.3}$ $23$M
JTT $\times$ $\surd$ $\times$ $86.7_{N/A}$ $93.3_{N/A}$ $23$M $81.1_{N/A}$ $88.0_{N/A}$ $23$M
PGI $\times$ $\surd$ $\times$ ${{79.5}}_{\pm 1.9}$ $95.5_{\pm 0.8}$ $23$M $85.3_{\pm 0.3}$ $87.3_{\pm 0.1}$ $23$M
CIM $\times$ $\surd$ $\times$ ${{77.2}}_{N/A}$ $95.6_{N/A}$ $23$M $83.6_{N/A}$ $90.6_{N/A}$ $23$M
CnC $\times$ $\surd$ $\times$ $88.5_{\pm 0.3}$ $90.9_{\pm 0.1}$ $23$M $88.9_{\pm 1.3}$ $88.9_{\pm 0.5}$ $23$M
DFR $\times$ $\surd$ $\times$ $92.9_{\pm 0.2}$ $94.2_{\pm 0.4}$ $23$M $88.3_{\pm 1.1}$ $91.3_{\pm 0.3}$ $23$M
UW $\surd$ $\surd$ $\surd$ $88.0_{\pm 1.3}$ $95.1_{\pm 0.3}$ $23$M $83.3_{\pm 2.8}$ $92.9_{\pm 0.2}$ $23$M
Subsample $\surd$ $\surd$ $\surd$ $86.9_{\pm 2.3}$ $89.2_{\pm 1.2}$ $23$M $86.1_{\pm 1.9}$ $91.3_{\pm 0.2}$ $23$M
LISA $\surd$ $\surd$ $\times$ $89.2_{\pm 0.6}$ $91.8_{\pm 0.3}$ $23$M $89.3_{\pm 1.1}$ $92.4_{\pm 0.4}$ $23$M
GroupDRO $\surd$ $\surd$ $\surd$ $86.7_{\pm 0.6}$ $93.2_{\pm 0.5}$ $23$M $86.3_{\pm 1.1}$ ${92.9}_{\pm 0.3}$ $23$M
CGD $\surd$ $\surd$ $\surd$ $88.9_{\pm 0.8}$ $91.3_{\pm 0.6}$ $23$M $90.0_{\pm 0.8}$ ${92.5}_{\pm 0.2}$ $23$M
DISC $\surd$ $\surd$ $\times$ ${{88.7}}_{\pm 0.4}$ $93.8_{\pm 0.7}$ $23$M N/A N/A N/A
PDE $\surd$ $\surd$ $\surd$ ${{90.3}}_{\pm 0.3}$ $92.4_{\pm 0.8}$ $23$M ${{91.0}}_{\pm 0.4}$ $92.0_{\pm 0.6}$ $23$M
CPT $\surd$ $\surd$ $\surd$ ${\mathbf{93.5}}_{\pm 0.4}$ ${{96.3}}_{\pm 0.2}$ $\mathbf{94}$k $\mathbf{92.0}_{\pm 0.3}$ ${93.2}_{\pm 0.3}$ $\mathbf{94}$k
De-bias ViT and CLIP Motivated by recent studies focusing on investigating spurious correlations of pre-trained transformer-based models [85, 18], we apply our proposed algorithm on ViT and CLIP and then compare them against those work. It is worth noting that, the authors did not experiment on CelebA with the blond hair target. Hence, we only show their original results on Waterbirds in Figure 3 and Table 2. As can be seen in Figure 3, while CPT enjoys much less computation and memory overhead (only $0.1\%$ and $0.04\%$ compared to ViT-B/16 itself and BiT-M-R50x3 [29]) it still outperforms the second-best model by a large margin. This significant improvement comes from our prompting design and optimization contribution, respectively. Details of backbones are given in Appendix B.
Figure 3: Results of fine-tuning ViT backbones on Waterbirds. Error bars represent the standard deviation over independent runs.
[…]
Table 2: Performance of CLIP models on Waterbirds. $\dagger$ indicates our implemented results. We use Bold font to indicate the best methods for different numbers of trainable parameters. N/A denotes those methods do not report the number of trainable parameters.
Model ResNet50 ViT-L/14@336px
Average Worst # Params Average Worst # Params
Pre-trained CLIP 90.890.8{90.8}90.8 $44.9$ 00 $88.5$ $34.0$ 00
Fine-tuned CLIP $81.3$ 77.177.1{77.1}77.1 $14789632$ 97.297.2{97.2}97.2 $89.7$ $786432$
ERM 93.593.5{93.5}93.5 $54.4$ $14789632$ $96.8$ $58.1$ $786432$
ERM Adapter $96.0$ 63.063.0{63.0}63.0 $524288$ $97.8$ $76.1$ $524288$
Contrastive Adapter $88.2$ 82.582.5{82.5}82.5 $263424$ $94.5$ $85.3$ $197632$
DFR $91.8$ $63.9$ N/A $96.1$ $65.9$ N/A
FairerCLIP $84.3$ $75.4$ N/A $92.2$ $86.0$ N/A
GroupDRO $83.3$ $73.7$ $14789632$ $94.1$ 90.890.8{90.8}90.8 $786432$
Yang et al. [85] $83.2$ 77.577.5{77.5}77.5 $14789632$ $96.9$ 90.590.5{90.5}90.5 $786432$
[…]
${}^{\dagger}$ 85.785.7{85.7}85.7 $64.6$ $8192$ $97.4$ $84.3$ $12288$
${}^{\dagger}$ $81.0$ $76.7$ $8192$ $95.8$ $90.5$ $12288$
CPT $83.3$ $78.4$ $8192$ $97.0$ $90.2$ $12288$
[…]
Table 3: Experimental results on MetaShift and ISIC dataset.
MetaShift ISIC
Method Average Worst AUROC # Params
ERM $72.9_{\pm 1.4}$ $62.1_{\pm 4.8}$ $36.4_{\pm 0.7}$ $23$M
UW $72.1_{\pm 0.9}$ $60.5_{\pm 3.8}$ $39.2_{\pm 0.6}$ $23$M
IRM $73.9_{\pm 0.8}$ $64.7_{\pm 2.1}$ $45.5_{\pm 3.6}$ $23$M
IB-IRM $74.8_{\pm 0.2}$ $65.6_{\pm 1.1}$ $38.6_{\pm 1.7}$ $23$M
V-REx $72.7_{\pm 1.7}$ $60.8_{\pm 5.5}$ $24.5_{\pm 6.4}$ $23$M
CORAL $73.6_{\pm 0.4}$ $62.8_{\pm 2.7}$ $37.9_{\pm 0.7}$ $23$M
Fish $64.4_{\pm 2.0}$ $53.2_{\pm 4.5}$ $42.0_{\pm 0.8}$ $23$M
GroupDRO $73.6_{\pm 2.1}$ $66.0_{\pm 3.8}$ $36.4_{\pm 0.9}$ $23$M
JTT $74.4_{\pm 0.6}$ $64.6_{\pm 2.3}$ $33.8_{\pm 0.0}$ $23$M
DM-ADA $74.0_{\pm 0.8}$ $65.7_{\pm 1.4}$ $35.8_{\pm 1.0}$ $23$M
LISA $70.0_{\pm 0.7}$ $59.8_{\pm 2.3}$ $38.0_{\pm 1.3}$ $23$M
DISC $75.5_{\pm 1.1}$ $73.5_{\pm 1.4}$ $55.1_{\pm 2.3}$ $23$M
CPT $\mathbf{79.8}_{\pm 2.6}$ $\mathbf{77.0}_{\pm{0.8}}$ $\mathbf{64.9}_{\pm 1.2}$ $\mathbf{47}$k
MetaShift and ISIC Results on those datasets are depicted in Table 3, along with the number of parameters required for each baseline. We mostly pick those that utilize group information during training for better benchmarking the efficiency of our proposed framework. In summary, the performance of CPT exceeds other methods by large margins, especially $\approx 10$ AUROC score on ISIC. On MetaShift, it is also the best method in terms of average and worst-group accuracy, improving the performance by approximately $4\%$.